OpenStudy (anonymous):

Solve for x: (x - 9)^2 = 1

OpenStudy (anonymous):

x - 9 = 1 x = 10

OpenStudy (anonymous):

\[x-9=1\] or \[x-9=-1\] \[x=10\] or \[x=8\]

OpenStudy (anonymous):

or x-9 = -1

OpenStudy (anonymous):

yeah 8 i forgot -1

OpenStudy (anonymous):

could you explain how you got this please

OpenStudy (anonymous):

you have to factor in the power in the problem